# Thermodynamic aspects of FLRW Universe in Einstein-Gauss-Bonnet domains

### v2

@@ I Introduction
The article is organized as follows: In Sect. II a general review of the main aspect of the Einstein–Gauss–Bonnet are provided. Besides, the equation of state for an FLRW metric is obtained from the thermodynamics unified first law, and a deep discussion about the final sign of the surface gravity and temperature is given, accompanied by an analysis of the causal nature of the apparent horizon and the possible matter distribution filling the Universe, next, in Sect. III, $P-v$ phase transitions are explored, providing a complete study about the incidence of the space-time dimensions on this phenomenon. Also, it provides a thorough assay of the “inverted” first-order phase transition undergone by the system in Sect. IV, the geometrothermodynamics and critical exponents are analyzed. For the former, we perform a detailed study about the sign of the normalized scalar curvature, accounting for phase transitions near the critical point. The critical exponents for the heat capacity at constant volume $C_{v}$, shear viscosity $\eta$, compressibility $\kappa_{T}$ and pressure $P$ are obtained. In Sect. V, we study the Duan’s topological approach for phase transitions and finally, in Sect. VI we conclude the present investigation.
@@ II.1 Field equations on a FLRW background
The full action of the $n$–dimensional Einstein-Gauss-Bonnet (EGB) gravity, coupled to matter fields, is given by
@@
where $\mathcal{R}$ is the $n$–dimensional Ricci scalar, B.T. stands for boundary terms in order to have a well-posed variational principle, $G_{n}$ the $n$–dimensional gravitational constant, $\alpha$ is a constant with units of length squared and $S_{m}$ is the action of the matter fields. The Gauss-Bonnet (GB) term $\mathcal{G}_{GB}$ reads as
@@
In the above expression $h_{ij}=\text{diag}\{-1,a^{2}/\left(1-kr^{2}\right)\}$ is the metric on $t-r$ plane (with $k=0,\pm 1$ is the spatial curvature), while $(n-2)$–dimensional unit sphere $d\Omega_{(n-2)}^{2}$ remains unaltered. Nevertheless, the function $R$, which is the areal radius, contains information of the $t-r$ chart, and it is defined as $R(t,r)\equiv a(t)r$.
@@
where $h^{ij}=\text{diag}\{-1,\left(1-kr^{2}\right)/a^{2}\}$ is the inverse of the metric $h_{ij}$. Solving (11) one obtains the following solution for the $AH$
@@ II.2 The equation of state
$\kappa_{\text{HK}}=2\pi\rho R_{AH}\left(\omega-\frac{1}{3}\right).$ (29)
@@
So, for inflation ($\omega=-1$), dust ($\omega=0$), and dark energy ($\omega=-1$), the surface gravity (29) is negative, while for radiation ($\omega=1/3$), it is zero (degenerated). On the other hand, for $\omega>1/3$, the surface gravity is positive. All the cases $\omega=\{-1,0,1/3\}$ correspond to the eras described by the $\Lambda$–CDM model, leading to an expanding cosmology. Regarding the temperature, it proportional to $T\propto+\kappa_{\text{HK}}$ for future $AH$ and $T\propto-\kappa_{\text{HK}}$ for past $AH$. From this information, one can extract the following combinations (Helou:2015yqa, ; Helou:2015zma, ): i) future/outer, ii) past/inner, iii) past/outer, and iv) future/inner horizons. The $AH$ is past/inner for expanding cosmologies and future/inner for contracting cosmologies. So, the final sign of the temperature in the former is positive and negative in the second case. Besides these scenarios, there is another one in the category of expanding cosmologies, where the matter distribution is the so-called stiff matter. In such a case, the barotropic EoS is given by $p(\rho)=\rho$, that is, with EoS parameter [^9] In table 1, we have placed for stiff matter the condition $\omega\geq 1$. Actually, the stiff matter takes the equality. However, in some speculative cosmological models, one can consider the EoS parameter greater than one. $\omega=1$. This kind of matter has been used in the study of the early Universe stage, leading to a classical bouncing model, free from singularities (Oliveira-Neto:2011uhf, ) and also to describe the same problem from the quantum point of view (Falciano:2007yf, ). For this type of matter distribution, the $AH$ is a past/outer one yielding to a negative temperature. Now, with this information at hand, we can be distinguished among three cases as summarized in table 1.
It should be noted that, in order to ensure the expanding scenario for the inflation and matter domination eras, the time variation of the $AH$ must be bounded from below. In the radiation era, the surface gravity degenerates; it is zero, leading to a null temperature. For the cosmological constant, the expansion can be considered as quasi-adiabatic or a very slow process, where the term proportional to $\dot{R}_{AH}$ is negligible.
@@
The type of matter distribution present in different epochs of the Universe’s evolution allows us to rapidly determine the sign of the surface gravity and then the temperature. Actually, this conclusion comes from $\gamma\equiv\mathcal{L}_{-}\theta_{-}/\mathcal{L}_{+}\theta_{-}$, that is, the ratio of Lie derivatives of the expansion coefficients. This determines the causal nature of a horizon (Dreyer:2002mx, ).
@@
Cosmological era Expanding/Contracting/Static Inner/Outer Future/Past Temperature Inflation ($\omega=-1$) Expanding ($\dot{R}_{AH}>0$) Inner ($\kappa_{\text{HK}}<0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow\frac{1}{2\pi R_{AH}}\left(1-\frac{\dot{R}_{AH}}{2HR_{AH}}\right)>0$, $10<\frac{\dot{R}_{AH}}{2HR_{AH}}<$ 1 Dust ($\omega=0$) Expanding ($\dot{R}_{AH}>0$) Inner ($\kappa_{\text{HK}}<0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow\frac{1}{2\pi R_{AH}}\left(1-\frac{\dot{R}_{AH}}{2HR_{AH}}\right)>0$, $10<\frac{\dot{R}_{AH}}{2HR_{AH}}<$ 1 Radiation ($\omega=1/3$) Expanding ($\dot{R}_{AH}>0$) Inner ($\kappa_{\text{HK}}=0$) Past $T\propto-\kappa_{\text{HK}}$ $\Rightarrow 0$ Cosmological constant ($\omega=-1$) Static ($\dot{R}_{AH}=0$) Inner ($\kappa_{\text{HK}}<0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow\frac{1}{2\pi R_{AH}}>0$ Stiff matter ($\omega\geq 1$) Expanding ($\dot{R}_{AH}>0$) Outer ($\kappa_{\text{HK}}>0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow\frac{1}{2\pi R_{AH}}\left(1-\frac{\dot{R}_{AH}}{2HR_{AH}}\right)<0$, $\frac{\dot{R}_{AH}}{2HR_{AH}}>1$ Phantom ($\omega\textless{}-1$) Contracting ($\dot{R}_{AH}<0$) Inner ($\kappa_{\text{HK}}<0$) Future $T\propto+\kappa_{\text{HK}}\Rightarrow-\frac{1}{2\pi R_{AH}}\left(1+\frac{|\dot{R}_{AH}|}{2HR_{AH}}\right)<0$
Table 1: The temperature expression (last column) for different types of cosmological eras. For inflation, matter, and stiff matter, the expression is the same, but the conditions on the signature of temperature are different, the latter having a negative temperature in contrast to the others. For the radiation case, the temperature is zero, while for the cosmological constant, there is no dependency on the variation of the $AH$. In this concern, the term proportional to $\dot{R}_{AH}$ can be discarded in the inflation era if one considers a super accelerated expansion, that is, $\frac{\dot{H}}{H^{2}}<<1\Rightarrow\ddot{a}(t)>0$ (deceleration parameter $q<0$).
@@
Now, the above simple analysis allows us to discriminate the expression for the temperature (last column in table 1) to be inserted in Eq. (24) according to the Universe era. Nevertheless, it should be taken into account that this analysis was done in order to give an interpretation of possible thermodynamic phenomena depending on the universe’s era. So, for a past/inner [^10] Except for the radiation and cosmological constant eras. and past/outer $AH$, the EoS (24) becomes [^11] It should be noted that for past/inner $AH$ the expression for the pressure is the same as given in Eq. (30).
@@ III P–V phase transitions
To explore the possibility of having $P-v$ phase transitions, the following conditions are required
@@
Figure 1: Left panel: The behavior of the pressure (30) against the molar volume for different values of the temperature. As can be observed, for those values above the critical one, the system is exhibiting an inverted first order phase transition (dashed blue line), where the wiggle now corresponds to stable states and the remaining one to unstable states. On the other hand, for temperature below the critical one, the system is completely unstable (dashed red line). Right panel: The reduced Gibbs free energy $\tilde{G}$ against the reduced pressure $\tilde{P}$. As the system undergoes a first-order phase transition, the shape of this grand potential is the classical swallowtail form, where the loop contains the unstable states of the system. For these plots we fixed $\alpha=0.2$ square length.
@@ IV.1 Ruppeiner Geometry
Figure 4: Top row: The normalized scalar curvature versus the volume. The left panel is showing the normalized curvature trend different temperatures. For some values, the scalar curvature is negative, positive, and again negative. This fluctuation reveals the presence of a phase transition. Points, where the scalar curvature nullifies, correspond to those values of ${T}$ satisfying ${R}_{N}=0$, giving rise to the so-called sign-changing curve. The right panel shows a zoom of the normalized curvature when it changes in sign for $T>T_{c}$. Bottom row: In this case, the normalized scalar curvature trend is quite different from the previous situation. Here, this invariant starts from a positive position for $T>T_{c}$ and then becomes negative in nature. This confirms that for $n=6$ no phase transitions are occurring. Besides, repulsive interactions (unstable states) and attractive interactions (stable states) exist. For these plots we fixed $\alpha=0.2$ square length.
@@ IV.2 Critical phenomena
As ${P}$ is linear in ${T}$, then there is no higher order in $\epsilon$. Now, the procedure dictates using Maxwell’s construction along with (49) to determine the small ${v}_{s}$ and large ${v}_{l}$ volumes in analytical way. Nevertheless, as this is not a genuine first order phase transition, and given that the unstable states are not confined, we cannot apply Maxwell’s construction to determine the small and large volumes at the coexistent locus. Therefore, it is not possible to estimate the behavior of the molar volume in terms of the reduced temperature $\epsilon$, namely, the critical exponent $\hat{\beta}$ cannot be determined for this cosmological model. Next, from the Eq. (49), retaining terms up to first one obtains
@@ VI Conclusions
  • The topological approach further corroborates the two previous analysis. In this concern, we found a generating point for $n=5$ accounting for the phase transition. In this case the system has a global topological charge equal to zero while for $n=6$, the scheme does not show any generating point, therefore no phase transitions are occurring. In this case it provides a global topological charge equal to one.
Despite the model does not show a realistic scenario, we can conclude that gravity theories formulated in dimensions higher than four, including high-order gravitational operators, offer a good scenario to test their feasibility (at least from the theoretical point of view) in the cosmological context by analyzing their thermodynamics description, microstructure and topological behavior. Of course, as in the present case, they can reveal intriguing and interesting phenomena. It is suggested to explore the same study reported here to higher order Lovelock terms, pure Lovelock gravity, or unified gravity theories such as Kaluza-Klein, for instance.

### v3

@@ I Introduction
The article is organized as follows: In Sect. II a general review of the main aspect of the Einstein–Gauss–Bonnet are provided. Besides, the equation of state for an FLRW metric is obtained from the thermodynamics unified first law, and a deep discussion about the final sign of the surface gravity and temperature is given, accompanied by an analysis of the causal nature of the apparent horizon and the possible matter distribution filling the Universe, next, in Sect. III, PvP-v phase transitions are explored, providing a complete study about the incidence of the space-time dimensions on this phenomenon. Also, it provides a thorough assay of the “inverted” first-order phase transition undergone by the system in Sect. IV, the geometrothermodynamics and critical exponents are analyzed. For the former, we perform a detailed study about the sign of the normalized scalar curvature, accounting for phase transitions near the critical point. The critical exponents for the heat capacity at constant volume $C_{v}$, shear viscosity $\eta$, compressibility $\kappa_{T}$ and pressure $P$ are obtained. In Sect. V, we study the Duan’s topological approach for phase transitions and finally, in Sect. VI we conclude the present investigation.
@@ II.1 Field equations on a FLRW background
The full action of the nn–dimensional Einstein-Gauss-Bonnet (EGB) gravity, coupled to matter fields, is given by
@@
where $\mathcal{R}$ is the nn–dimensional Ricci scalar, B.T. stands for boundary terms in order to have a well-posed variational principle, $G_{n}$ the nn–dimensional gravitational constant, $\alpha$ is a constant with units of length squared and $S_{m}$ is the action of the matter fields. The Gauss-Bonnet (GB) term $\mathcal{G}_{GB}$ reads as
@@
In the above expression $h_{ij}=\text{diag}\{-1,a^{2}/\left(1-kr^{2}\right)\}$ is the metric on trt-r plane (with $k=0,\pm 1$ is the spatial curvature), while (n2)(n-2)–dimensional unit sphere $d\Omega_{(n-2)}^{2}$ remains unaltered. Nevertheless, the function $R$, which is the areal radius, contains information of the trt-r chart, and it is defined as $R(t,r)\equiv a(t)r$.
@@
where $h^{ij}=\text{diag}\{-1,\left(1-kr^{2}\right)/a^{2}\}$ is the inverse of the metric $h_{ij}$. Solving (11) one obtains the following solution for the AHAH
@@ II.2 The equation of state
$\kappa_{\text{HK}}=\frac{4\pi G_{n}\rho}{(n-2)}\frac{(\omega+1)(n-1)-4}{(n-1)}R_{AH}+\frac{\tilde{\alpha}}{R^{3}_{AH}}(1-\dot{R}_{AH}).$ (29)
@@
As can be seen, the signature of the surface gravity depends on several factors: i) The EoS parameter, ii) the dimensions $n$ and iii) the GB corrections through the rate expansion/contraction of the Universe. In this concern, as the temperature is proportional to the surface gravity, that is, $T\propto\kappa_{\text{HK}}$, it is so important to determine whether $\kappa_{\text{HK}}$ is positive or negative in nature. Particularly, one has $T\propto+\kappa_{\text{HK}}$ for future $AH$ and $T\propto-\kappa_{\text{HK}}$ for past $AH$. The future or past character of the $AH$ depends on the cosmological scenario, that is, contracting or expanding, respectively. While the inner/outer feature concerns the sign of the surface gravity, being inner for negative and outer for positive surface gravity. In general, one has the following combinations (Helou:2015yqa, ; Helou:2015zma, ): i) future/outer, ii) past/inner, iii) past/outer, and iv) future/inner horizons. The $AH$ is past/inner for expanding cosmologies and future/inner for contracting cosmologies. So, the final sign of the temperature in the former is positive and negative in the second case. Besides these scenarios, there is another one in the category of expanding cosmologies, where the matter distribution is the so-called stiff matter. In such a case, the barotropic EoS is given by $p(\rho)=\rho$, that is, with EoS parameter [^9] In table 1, we have placed for stiff matter the condition $\omega\geq 1$. Actually, the stiff matter takes the equality. However, in some speculative cosmological models, one can consider the EoS parameter greater than one. $\omega=1$. This kind of matter has been used in the study of the early Universe stage, leading to a classical bouncing model, free from singularities (Oliveira-Neto:2011uhf, ) and also to describe the same problem from the quantum point of view (Falciano:2007yf, ).
Now, if one takes into account the usual matter distributions whose have dominated the different epochs, that is, inflation ($\omega=-1$), dust ($\omega=0$), dark energy ($\omega=-1$) and radiation ($\omega=1/3$), all of them correspond to inner/past $AH$ with a positive temperature. However, in the GR case, the radiation epoch has a degenerate $AH$, that is, an $AH$ where $\kappa_{\text{HK}}=0$ consequently $T=0$ (Helou:2015yqa, ; Helou:2015zma, ). Nevertheless, when GB corrections are present, this is no longer the case. Furthermore, for the inflation and cosmological constant scenario, one needs to demand $\dot{R}_{AH}=0$, that is, a quasi-adiabatic process. On the other hand, dust epoch and radiation entail $\dot{R}_{AH}>1$ and $\dot{R}_{AH}>>1$, respectively. The era of hypothetic stiff matter imposes $\dot{R}_{AH}>0$ leading to $T<0$. Finally, in the event of a contracting scenario due to the distribution of phantom matter, the $AH$ is inner/future in nature with $T<0$. For this to happen, one needs $\dot{R}_{AH}<<0$.
@@
In summary, the type of matter distribution present in different epochs of the Universe’s evolution allows us to rapidly determine the sign of the surface gravity and then the temperature. Actually, this conclusion comes from $\gamma\equiv\mathcal{L}_{-}\theta_{-}/\mathcal{L}_{+}\theta_{-}$, that is, the ratio of Lie derivatives of the expansion coefficients. This determines the causal nature of a horizon (Dreyer:2002mx, ). Table 1 shows some of the possibilities given in the GB context.
@@
Cosmological era Expanding/Contracting Inner/Outer Future/Past Temperature Inflation ($\omega=-1$) Expansion ($\dot{R}_{AH}=0$) Inner ($\kappa_{\text{HK}}<0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow T>0$ Dust ($\omega=0$) Expansion ($\dot{R}_{AH}>0$) Inner ($\kappa_{\text{HK}}<0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow T>0$ Radiation ($\omega=1/3$) Expansion ($\dot{R}_{AH}>0$) Inner ($\kappa_{\text{HK}}<0$) Past $T\propto-\kappa_{\text{HK}}$ $\Rightarrow T>0$ Cosmological constant ($\omega=-1$) Expansion ($\dot{R}_{AH}=0$) Inner ($\kappa_{\text{HK}}<0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow T>0$ Stiff matter ($\omega\geq 1$) Expansion ($\dot{R}_{AH}>0$) Outer ($\kappa_{\text{HK}}>0$) Past $T\propto-\kappa_{\text{HK}}\Rightarrow T<0$ Phantom ($\omega\textless{}-1$) Contracting ($\dot{R}_{AH}<0$) Inner ($\kappa_{\text{HK}}<0$) Future $T\propto+\kappa_{\text{HK}}\Rightarrow T<0$
Table 1: The temperature expression (last column) for different types of cosmological eras.
@@
Now, the above simple analysis allows us to discriminate the expression for the temperature (last column in table 1) to be inserted in Eq. (24) according to the Universe era. Nevertheless, it should be taken into account that this analysis was done in order to give an interpretation of possible thermodynamic phenomena depending on the universe’s era. So, for a past/inner and past/outer $AH$, the EoS (24) becomes
@@ III P–V phase transitions
To explore the possibility of having PvP-v phase transitions, the following conditions are required
@@
Figure 1: Left panel: The behavior of the pressure (30) against the molar volume for different values of the temperature. As can be observed, for those values above the critical one, the system is exhibiting an inverted first order phase transition (dashed blue line), where the wiggle now corresponds to stable states and the remaining one to unstable states. On the other hand, for temperature below the critical one, the system is completely unstable (dashed red line). Right panel: The pressure and the “spinodal” curve versus the volume. For these plots we fixed $\alpha=0.2$ square length.
@@ IV.1 Ruppeiner Geometry
Figure 4: Top row: The normalized scalar curvature versus the volume. The left panel is showing the normalized curvature trend different temperatures. For some values, the scalar curvature is negative, positive, and again negative. This fluctuation reveals the presence of a phase transition. Points, where the scalar curvature nullifies, correspond to those values of T{T} satisfying ${R}_{N}=0$, giving rise to the so-called sign-changing curve. The right panel shows a zoom of the normalized curvature when it changes in sign for $T>T_{c}$. Bottom row: In this case, the normalized scalar curvature trend is quite different from the previous situation. Here, this invariant starts from a positive position for $T>T_{c}$ and then becomes negative in nature. This confirms that for $n=6$ no phase transitions are occurring. Besides, repulsive interactions (unstable states) and attractive interactions (stable states) exist. For these plots we fixed $\alpha=0.2$ square length.
@@ IV.2 Critical phenomena
As P{P} is linear in T{T}, then there is no higher order in $\epsilon$. Now, the procedure dictates using Maxwell’s construction along with (49) to determine the small ${v}_{s}$ and large ${v}_{l}$ volumes in analytical way. Nevertheless, as this is not a genuine first order phase transition, and given that the unstable states are not confined, we cannot apply Maxwell’s construction to determine the small and large volumes at the coexistent locus. Therefore, it is not possible to estimate the behavior of the molar volume in terms of the reduced temperature $\epsilon$, namely, the critical exponent $\hat{\beta}$ cannot be determined for this cosmological model. Next, from the Eq. (49), retaining terms up to first one obtains
@@ VI Conclusions
  • The topological approach further corroborates the two previous analysis. In this concern, we find a generating point for $n=5$ that accounts for the phase transition. In this case, the system has a global topological charge equal to zero while for $n=6$, the scheme does not show any generating point; therefore, no phase transitions are occurring. In this case, it provides a global topological charge equal to one.
Despite the fact that the model does not show a realistic scenario, we can conclude that gravity theories formulated in dimensions higher than four, including high-order gravitational operators, offer a good scenario to test their feasibility (at least from the theoretical point of view) in the cosmological context by analyzing their thermodynamics description, microstructure, and topological behavior. Of course, as in the present case, they can reveal intriguing and interesting phenomena. It is suggested to explore the same study reported here for higher-order Lovelock terms, pure Lovelock gravity, or unified gravity theories such as Kaluza-Klein, for instance. Additionally, including a scalar field action for the action is an interesting issue for EGB gravity. The scalar field profoundly impacts the theory under consideration; for example, in four dimensions, it is well known that the EGB term is a topological invariant term and does not contribute to cosmological dynamics through the equation of motion. Including the scalar field dramatically changes this situation; in this case, the EGB term does contribute to the dynamics with a rich non-linear structure, and therefore, in an FLRW universe, it could be possible to find phase transition as we describe in the present article. Also, we need to study what kind of scalar function we are considering in higher dimensions to obtain the cosmological behavior and, therefore, the subsequent thermal history of the universe where, of course, we hope to modify our results on phase transition. These modifications are mainly supported because there is a strong suggestion about the relationship between the first law of thermodynamics of the apparent horizon and the Friedmann equation. To our knowledge, a critical point is that no one has studied the apparent thermodynamics for the case where the geometry is coupled with a single scalar field. We left the interesting model with the EGB term and the inclusion of scalar field coupling as a future work. Also, we did not address the issue of examining the tensor perturbation EGB theories in light of the GW170817 results, as discussed in the references Oikonomou:2024etl , Oikonomou:2021kql , Oikonomou:2025ccs and Odintsov:2025kyw . Those results and their effect on the phase transition pattern could be essential for testing our model and its theoretical predictions, and we will discuss those points elsewhere.